# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@ 2 Four Definitions of Free Sets
Figure 1: Showing that every proper-good set is a collinear set. Dummy vertices are hollow circles and elements of $S$ are filled circles.
They then use Tutte’s Convex Embedding Theorem [57] twice: Once to find a straight-line crossing-free drawing of $G^{\prime}[X\cup Y]$ with all vertices of $X$ below the ${x}$-axis and all vertices of $Y$ on the ${x}$-axis, and a second time to find a straight-line crossing-free drawing of $G^{\prime}[Y\cup Z]$ with the vertices of $Z$ above the ${x}$-axis. This gives a straight-line crossing-free drawing $\Gamma^{\prime}$ of $G^{\prime}$, as depicted in Fig. 1(b). The straight-line crossing-free drawing $\Gamma^{\prime}$ is almost a straight-line crossing-free drawing of $G$ except that the edges that properly intersect $C$ are represented by two line segments, one below and one above the ${x}$-axis. The authors then apply a result of Pach and Tóth [47] that shows the edges in this crossing-free drawing can be straightened without changing the ${y}$-coordinate of any vertex, as depicted in Fig. 1(c). In particular, in the resulting straight-line crossing-free drawing of $G$, the vertices of $G$ intersected by $C$ remain on the ${x}$-axis, as required.
A few notes on the proof are in order. First, we recall that the proof of Tutte’s Convex Embedding Theorem is mostly algebraic: it involves setting up a system of linear equations whose solution determines the coordinates of the vertices in the straight-line crossing-free drawing. Second, we note that the preceding proof almost proves that every proper-good set is a free-collinear set. Tutte’s theorem allows us to choose the ${x}$-coordinates of the vertices in $Y$, including the vertices in $S$. However, the straightening step, which converts $\Gamma^{\prime}$ into $\Gamma$ by straightening the edges that cross the ${x}$-axis, changes the ${x}$-coordinates of the vertices.
@@
The proof that every collinear set is a free-collinear set is due to Dujmović et al. [29]. Prior to this, Da Lozzo et al. [24] had established the equivalence of collinear and free-collinear sets for the class of planar $3$-trees and asked again if this equivalence was true in general. In subsequent years this repeatedly lead to the question (originally posed by Ravsky and Verbitsky [48]) “is every collinear set free?”. The proof of Dujmović et al. [29] has two main parts, which we outline below. If $S:=(v_{1},\ldots,v_{s})$ is an ordered collinear set, then $G$ has a straight-line crossing-free drawing $\Gamma$ in which the vertices of $S$ appear, in order, on the ${x}$-axis. The goal is to show that, for any $x_{1}<\cdots<x_{s}$, the vertex $v_{i}$ can be moved to $(x_{i},0)$, for each $i\in\{1,\ldots,s\}$, such that the resulting drawing is a straight-line crossing-free drawing. By adding (straight-line) edges to $\Gamma$, we may assume that $G$ is a triangulation. In the following, when we say that an edge or a face properly intersects the ${x}$-axis, we mean that the edge or the face contains points that are strictly above and points that are strictly below the ${x}$-axis.
The first part of the proof is combinatorial. In this part, a sequence of modifications is done to $G$ with the end goal of obtaining a triangulation $G^{\prime}$ that has an (no longer straight-line) crossing-free drawing $\Gamma^{\prime}$ with following properties (see Fig. 2(c)):
@@
    For each vertex $v\in S$ there are exactly two faces of $G^{\prime}$ incident on $v$ that properly intersect the ${x}$-axis.
@@
To obtained $G^{\prime}$ we apply three types of operations, two of which are illustrated in Fig. 2. (1) If some edge with no endpoint in $S$ is incident on two faces, neither of which properly intersect the ${x}$-axis, then this edge is contracted (the orange edge in Fig. 2(a) meets those conditions and its contraction results in an embedded graph depicted in Fig. 2(b)). This operation eliminates these two faces. (2) If some edge with no endpoints in $S$ is incident on one face whose closure is disjoint from the ${x}$-axis, whose other incident face properly intersects the ${x}$-axis, and meets some other technical conditions, [^5] These technical conditions ensure that, after drawing $G^{\prime}$ the union of the two faces incident on the flipped edge is a convex quadrilateral. then this edge is flipped (the lilac edge in Fig. 2(b) meets the necessary conditions and its flipping results in an embedded graph depicted in Fig. 2(c)). This edge flip replaces the two faces incident to the original edge with two faces that each properly intersect the ${x}$-axis. (3) Finally, any edge with both endpoints in $S$ is flipped. After this flip, the flipped edge and the two faces incident to it properly intersect the ${x}$-axis. These operations (contractions and flips) are done exhaustively, one by one, in any order. The operations maintain that $S$ is a proper-good set in $G^{\prime}$. As discussed above, either of these operations may introduce a separating triangle in $G^{\prime}$. How this is handled is explained below. Once these operations are done exhaustively, removing the edges of $G^{\prime}$ both of whose endpoints are either strictly above or strictly below ${x}$-axis (see the green edges in Fig. 2(c)) yields a structure, called an A-graph (see Fig. 2(d)), in which every face is either a quadrilateral whose four edges each intersect the ${x}$-axis or have an endpoint on the ${x}$-axis or a triangle with vertices above, on, and below the ${x}$-axis. The main task now is to find a straight-line crossing-free drawing of this $A$-graph so that the vertices of $S$ are moved to the specified locations on the ${x}$-axis. Once we achieve that, the edge-removal, flipping, and contracting operations are easily undone without changing the locations of vertices in $S$, resulting in straight-line crossing-free drawing of $G$ with the vertices of $S$ at their specified locations on the ${x}$-axis.
@@
Finding a straight-line crossing-free drawingof this $A$-graph so that the vertices of $S$ are placed to the specified locations on the ${x}$-axis is the goal of the second part of the proof, which involves a combination of linear algebra and graph theory. The authors prove a very strong statement about the A-graph $G^{\prime}$. They show that it is possible to obtain a straight-line crossing-free drawing of $G^{\prime}$ in which the intersection point of the closure of each edge with the ${x}$-axis is specified, subject to the ordering constraints imposed by the crossing-free drawing of $G^{\prime}$. Note that this implies that $S$ is a free-collinear set in $G^{\prime}$ since each vertex in $S$ is incident to an edge of $G^{\prime}$ (recall that no edge incident to a vertex of $S$ is flipped or contracted unless both of its endpoints are on the ${x}$-axis). The vertex $v_{i}$ has at least three incident edges in $G^{\prime}$. By specifying that the closure of each of these edges must intersect the ${x}$-axis at $(x_{i},0)$ we ensure that $v_{i}$ is mapped to $(x_{i},0)$.
@@
In addition to specifying the location where each edge crosses the ${x}$-axis, the locations of the three vertices $x,y,z$ on the convex hull of this crossing-free drawing (the vertices marked with   in Fig. 2(d)) can be chosen, provided that the choice agrees with the choice of intersection points chosen for the edges induced by $x,y,z$ (the intersections marked with $\times$ in Fig. 2(d)). This latter requirement justifies the assumption that the graph does not contain separating triangles: The subgraph in the closure of the exterior of a separating triangle $xyz$ can be drawn inductively. This determines the locations of $x$, $y$, and $z$ which are then specified when inductively drawn the graph in the closure of the interior of the separating triangle.
@@
The proof works by assigning slopes to the edges in $G^{\prime}$. To avoid vertical edges, it is helpful to rotate the coordinate system by $90$ degrees so that the roles of the x- and y-axes swap. To prove this stronger claim about A-graphs, the authors set up a system $M$ of linear equations whose variables are the slopes of the edges of $G^{\prime}$ and whose coefficients are determined by the (given) intersection location of each edge with the ${y}$-axis. Most of the equations in $M$ correspond to the fact that three or more edges incident to a common vertex $v$ must intersect in a single point (the location of $v$). [^6] Some additional constraints, call proportionality constraints, are placed on the slopes of edges incident to vertices on the ${y}$-axis.
@@
The hardest step in this proof is to show that the linear system $M$ has a unique solution. Here, the authors leverage the fact that $S$ is a proper-good set in $G^{\prime}$ and, therefore, a collinear set in $G^{\prime}$. Therefore, $G^{\prime}$ has a straight-line crossing-free drawing $\Gamma_{0}$ in which the vertices of $S$ appear, in order, on the ${y}$-axis. From this straight-line crossing-free drawing $\Gamma_{0}$, one can read off a system $M_{0}$ of linear equations that has the same structure as $M$ but with different coefficients (that are determined by the ${y}$-coordinates of vertices in $S$ in $\Gamma_{0}$). The system $M_{0}$ has a solution, namely the solution given by the slopes of the edges in $\Gamma_{0}$. The authors first show that this solution to $M_{0}$ is unique. The final step involves continuously modifying the coefficients in $M_{0}$ to obtain a continuum of linear systems $M_{t}$, $0\leq t\leq 1$ where $M_{1}=M$. Using the fact that $M_{0}$ has a unique solution as a starting point, they are then able to show that $M_{t}$ has a unique solution for each $0\leq t\leq 1$. In particular, $M_{1}=M$ has a solution, which determines the desired straight-line crossing-free drawing of $G^{\prime}$
@@
The equivalence between free-collinear sets and free sets has a disappointingly easy proof. The following argument appears in Bose et al. [17], but similar “perturb and scale” arguments are fairly common.
Let $S:=(v_{1},\ldots,v_{s})$ be a free-collinear set in $G$ and let $x_{1}<\cdots<x_{s}$ and $y_{1},\ldots,y_{s}$ be real numbers. Since $S$ is a free-collinear set in $G$, $G$ has a straight-line crossing-free drawing in $\Gamma_{0}$ where $v_{i}$ is at $(x_{i},0)$ for each $i\in\{1,\ldots,s\}$. It follows easily from the definition of straight-line crossing-free drawing that there exists some $\epsilon>0$ such that perturbing each of the vertices in $\Gamma_{0}$ by at most $\epsilon$ results in another straight-line crossing-free drawing $\Gamma_{0}^{\prime}$. Let $y:=\max\{|y_{1}|,\ldots,|y_{s}|\}$. Define $\Gamma_{0}^{\prime}$ by moving $v_{i}$ to $(x_{i},\epsilon y_{i}/y)$, for each $i\in\{1,\ldots,s\}$ and leaving the other vertices of $G$ fixed. Then each vertex in $\Gamma_{0}^{\prime}$ has been moved a distance of at most $\epsilon$ from its location in $\Gamma_{0}$, so $\Gamma_{0}^{\prime}$ is a straight-line crossing-free drawing of $G$. Finally, multiply the ${y}$-coordinate of each vertex by $y/\epsilon$ to obtain a final straight-line crossing-free drawing $\Gamma$. This y-scaling is an affine transformation that does not introduce any crossings, so $\Gamma$ is a straight-line crossing-free drawing of $G$ and, for each $i\in\{1,\ldots,s\}$, $\Gamma$ places $v_{i}$ at $(x_{i},y_{i})$.
@@ 3 Graph Classes with Large Free Sets
What is the maximum size of a free set that is guaranteed to exist in any $n$-vertex planar graph? Ravsky and Verbitsky [48] observed that upper bounds on this value can be obtained from existing work on the circumference of cubic triconnected planar graphs. They observe that if a triangulation $G$ has a collinear set of size $\ell$ (and thus a large free set by 1), then its dual graph $G^{*}$ (which is a cubic triconnected planar graph) has a cycle of length $\Omega(\ell)$. The length $c(G^{*})$ of a longest cycle in $G^{*}$ is called the circumference of $G^{*}$. The circumference of cubic triconnected planar graphs has a long and rich history dating back to at least 1884 when Tait [55] conjectured that every such graph is Hamiltonian.
Tait’s Conjecture was famously disproved in 1946 by Tutte, who gave an example of a non-Hamiltonian cubic triconnected planar graph having 46 vertices [56]. Repeatedly replacing vertices of Tutte’s graph with copies of itself gives a family of graphs, $\langle G_{i}:i\in\mathbb{N}\rangle$ in which $G_{i}$ has $46\cdot 45^{i}$ vertices and circumference at most $45\cdot 44^{i}$. Stated another way, $n$-vertex members of the family have circumference $O(n^{\sigma})$, for $\sigma=\log_{45}(44)<0.9941$. The current best upper bound of this type is due to Grünbaum and Walther [40] who construct a family of cubic triconnected planar graphs in which $n$-vertex members have circumference $O(n^{\sigma})$ for $\sigma=\log_{23}(22)<0.9859$. The dual of such a graph is a triangulation having $\Theta(n)$ vertices whose largest free set has size $O(n^{\sigma})$.
@@
Owens [45] constructs a family of $n$-vertex cubic triconnected planar graphs whose faces have size at most $7$, and that contain no cycle of length $\Omega(n^{0.9976})$. The dual of such a graph is a triangulation having $\Theta(n)$ vertices and maximum-degree $7$ whose largest free set has size $O(n^{0.9976})$.
@@
Cano et al. [20] also use upper bounds on the circumference of cubic triconnected planar graphs, along with the Erdős-Szekeres Theorem to show that $\operatorname{fix}_{\mathcal{G}}(n)\in O(n^{0.4948})$.
Since not all planar graphs have linear size free sets, we first consider the subclasses of planar graphs that do admit such large free sets (in Section 3.1). We then study general planar graphs (in Section 3.2). The relationship between these two key graph parameters (largest free set and circumference) is then used to obtain strong bounds for bounded-degree planar graphs (in Section 3.3).
3.1 Subclasses of planar graphs with linear-sized free sets
A level planar drawing of a graph is a straight-line crossing-free drawing  in the plane, such that the vertices are placed on a sequence of parallel lines (called levels), where each edge joins vertices in two consecutive levels. For example, a natural straight-line crossing-free drawing of a tree $T$, with root $r$, places each vertex of $T$ that is at distance $i$ from $r$ on the line $y=-i$. Thus, trees have level planar drawings.
The following easy argument shows that $n$-vertex graphs that have level planar drawings have free sets of size at least $n/2$. See Fig. 3. Starting with a level planar drawing $\Gamma$, create a simple closed curve $C$ that coincides with the drawing in some set of levels. If $C$ does not contain any two consecutive levels of $\Gamma$, then $C$ intersects $\Gamma$ only in the vertices of $G$ that form an independent set. Such a curve is thus a proper-good curve, and by having it intersect every second level of $G$ one can ensure that $\Gamma$ contains at least $\lceil n/2\rceil$ vertices of $G$.

### v2

@@ 2 Four Definitions of Free Sets
Figure 1: Showing that every proper-good set is a collinear set. Dummy vertices are hollow circles, and elements of $S$ are filled circles.
The proof then uses Tutte’s Convex Embedding Theorem [58] twice: Once to find a straight-line crossing-free drawing of $G^{\prime}[X\cup Y]$ with all vertices of $X$ below the ${x}$-axis and all vertices of $Y$ on the ${x}$-axis, and a second time to find a straight-line crossing-free drawing of $G^{\prime}[Y\cup Z]$ with the vertices of $Z$ above the ${x}$-axis. This gives a straight-line crossing-free drawing $\Gamma^{\prime}$ of $G^{\prime}$, as depicted in Fig. 1(b). More precisely, the proof applies a variant of Tutte’s Convex Embedding Theorem to a supergraph of $G^{\prime\prime}$ of $G^{\prime}[X\cup Y]$ (respectively, $G^{\prime}[Y\cup Z]$) whose outer face consists of a cycle that contains all the vertices of $Y$ plus a newly-introduced vertex $v$ [^7] Tutte’s Convex Embedding Theorem, as it appears in [58] requires that the vertices of the outer face be mapped to a strictly convex polygon. However, the proof holds in the current setting because (i) the vertices of the outer face of $G^{\prime\prime}$ induce a cycle and (ii) for each vertex $v$ of $G^{\prime}[X\cup Y]$ not on the outer face, $N_{G}(v)$ does not consist entirely of collinear vertices on the outer face. (The extra vertices and edges of these supergraphs are grey in Fig. 1(b).)
The straight-line crossing-free drawing $\Gamma^{\prime}$ is almost a straight-line crossing-free drawing of $G$ except that the edges that properly intersect $C$ are represented by two line segments, one below and one above the ${x}$-axis. The proof then applies a result of Pach and Tóth [49] (Theorem 1.2 in [49]) that shows the edges in this crossing-free drawing can be straightened without changing the ${y}$-coordinate of any vertex, as depicted in Fig. 1(c). In particular, in the resulting straight-line crossing-free drawing of $G$, the vertices of $G$ intersected by $C$ remain on the ${x}$-axis, as required.
A few notes on the proof are in order. First, we recall that the proof of Tutte’s Convex Embedding Theorem is mostly algebraic: it involves setting up a system of linear equations whose solution determines the coordinates of the vertices in the straight-line crossing-free drawing. The proof that every collinear set is a free-collinear set, sketched below, has a similar algebraic part. Second, we note that the preceding proof almost proves that every proper-good set is a free-collinear set. Tutte’s theorem allows us to choose the ${x}$-coordinates of the vertices in $Y$, including the vertices in $S$. However, the straightening step, which converts $\Gamma^{\prime}$ into $\Gamma$ by straightening the edges that cross the ${x}$-axis, changes the ${x}$-coordinates of the vertices.
@@
The proof that every collinear set is a free-collinear set is due to Dujmović et al. [28]. Prior to this, Da Lozzo et al. [24] had established the equivalence of collinear and free-collinear sets for the class of planar $3$-trees and asked again if this equivalence was true in general. In subsequent years, this repeatedly led to the question (originally posed by Ravsky and Verbitsky [50]) “is every collinear set free?”. The proof of Dujmović et al. [28] has two main parts, which we outline below.
If $S:=(v_{1},\ldots,v_{s})$ is a collinear set in $G$ then it is also a proper-good set in $G$. By adding one vertex in the outer face of $G$ and adding (not-necessarily straight-line) edges, we may assume that $G$ is a triangulation (and $S$ is still a proper-good set). As discussed in the previous section, the fact that $S$ is a proper-good set implies that $G$ is a collinear set, so $G$ has a straight-line crossing-free drawing $\Gamma$ in which the vertices of $S$ appear, in order, on the ${x}$-axis. The goal is to show that, for any $x_{1}<\cdots<x_{s}$, the vertex $v_{i}$ can be moved to $(x_{i},0)$, for each $i\in\{1,\ldots,s\}$, such that the resulting drawing is a straight-line crossing-free drawing. In the following, when we say that an edge or a face properly intersects the ${x}$-axis, we mean that the edge or the face contains points that are strictly above and points that are strictly below the ${x}$-axis.
The first part of the proof is combinatorial. In this part, a sequence of modifications is done to $G$ with the end goal of obtaining a triangulation $G^{\prime}$ that has a (no longer straight-line) crossing-free drawing $\Gamma^{\prime}$ with the following properties (see Fig. 2(c)):
@@
    For each vertex $v\in S$ there are exactly two faces of $G^{\prime}$ incident to $v$ that properly intersect the ${x}$-axis.
@@
To obtain $G^{\prime}$, we apply three types of operations, two of which are illustrated in Fig. 2. (1) If some edge with no endpoint in $S$ is incident on two faces, neither of which properly intersect the ${x}$-axis, then this edge is contracted (the orange edge in Fig. 2(a) meets those conditions and its contraction results in an embedded graph depicted in Fig. 2(b)). This operation eliminates these two faces. (2) If some edge with no endpoints in $S$ is incident on one face whose closure is disjoint from the ${x}$-axis, whose other incident face properly intersects the ${x}$-axis, and meets some other technical conditions, [^8] These technical conditions ensure that, after drawing $G^{\prime}$ the union of the two faces incident on the flipped edge is a convex quadrilateral. then this edge is flipped (the lilac edge in Fig. 2(b) meets the necessary conditions and its flipping results in an embedded graph depicted in Fig. 2(c)). This edge flip replaces the two faces incident to the original edge with two faces that each properly intersect the ${x}$-axis. [^9] The assumption that $G$ has no separating triangles ensures that this flipping operation does not introduce parallel edges, i.e., each edge of $G$ is flippable. (3) Finally, any edge with both endpoints in $S$ is flipped. After this flip, the flipped edge and the two faces incident to it properly intersect the ${x}$-axis. These operations (contractions and flips) are done exhaustively, one by one, in any order. The operations maintain that $S$ is a proper-good set in $G^{\prime}$. As discussed above, either of these operations may introduce a separating triangle in $G^{\prime}$. How this is handled is explained below. Once these operations are done exhaustively, removing the edges of $G^{\prime}$ both of whose endpoints are either strictly above or strictly below ${x}$-axis (see the green edges in Fig. 2(c)) yields a structure, called an A-graph (see Fig. 2(d)), in which every face is either a quadrilateral whose four edges each intersect the ${x}$-axis or have an endpoint on the ${x}$-axis or a triangle with vertices above, on, and below the ${x}$-axis. The main task now is to find a straight-line crossing-free drawing of this $A$-graph so that the vertices of $S$ are moved to the specified locations on the ${x}$-axis. Once we achieve that, the edge-removal, flipping, and contracting operations are easily undone without changing the locations of vertices in $S$, resulting in straight-line crossing-free drawing of $G$ with the vertices of $S$ at their specified locations on the ${x}$-axis.
@@
Finding a straight-line crossing-free drawing of this $A$-graph so that the vertices of $S$ are placed at the specified locations on the ${x}$-axis is the goal of the second part of the proof, which involves a combination of linear algebra and graph theory. The proof establishes a very strong statement about the A-graph $G^{\prime}$. The proof shows that it is possible to obtain a straight-line crossing-free drawing of $G^{\prime}$ in which the intersection point of the closure of each edge with the ${x}$-axis is specified, subject to the ordering constraints imposed by the crossing-free drawing of $G^{\prime}$. Note that this implies that $S$ is a free-collinear set in $G^{\prime}$ since each vertex in $S$ is incident to an edge of $G^{\prime}$ (recall that no edge incident to a vertex of $S$ is ever contracted and no edge incident to a vertex of $S$ is flipped unless both of its endpoints are on the ${x}$-axis). Each vertex $v_{i}\in S$ has at least three incident edges in $G^{\prime}$. By specifying that the closure of each of these edges must intersect the ${x}$-axis at $(x_{i},0)$, we ensure that $v_{i}$ is mapped to $(x_{i},0)$.
@@
In addition to specifying the location where each edge crosses the ${x}$-axis, the locations of the three vertices $x,y,z$ on the convex hull of this crossing-free drawing (the vertices marked with   in Fig. 2(d)) can be chosen, provided that the choice agrees with the choice of intersection points chosen for the edges induced by $x,y,z$ (the intersections marked with $\times$ in Fig. 2(d)). This latter requirement justifies the assumption that the graph does not contain separating triangles: The subgraph in the closure of the exterior of a separating triangle $xyz$ can be drawn inductively. This determines the locations of $x$, $y$, and $z$, which are then specified when inductively drawing the graph in the closure of the interior of the separating triangle.
@@
The proof works by assigning slopes to the edges in $G^{\prime}$. To avoid vertical edges, it is helpful to rotate the coordinate system by $90$ degrees so that the roles of the ${x}$- and ${y}$-axes swap. To prove this stronger claim about A-graphs, the proof sets up a system $M$ of linear equations whose variables are the slopes of the edges of $G^{\prime}$ and whose coefficients are determined by the (given) intersection location of each edge with the ${y}$-axis. Most of the equations in $M$ correspond to the fact that three or more edges incident to a common vertex $v$ must intersect in a single point (the location of $v$). [^10] Some additional constraints, called proportionality constraints, are placed on the slopes of edges incident to vertices on the ${y}$-axis.
@@
The hardest step in this proof is to show that the linear system $M$ has a (unique) solution. Here, the proof leverages the fact that $S$ is a proper-good set in $G^{\prime}$ and, therefore, a collinear set in $G^{\prime}$. Therefore, $G^{\prime}$ has a straight-line crossing-free drawing $\Gamma_{0}$ in which the vertices of $S$ appear, in order, on the ${y}$-axis. From this straight-line crossing-free drawing $\Gamma_{0}$, one can read off a system $M_{0}$ of linear equations that has the same structure as $M$ but with different coefficients (that are determined by the ${y}$-coordinates of vertices in $S$ in $\Gamma_{0}$). The system $M_{0}$ has a solution, namely the solution given by the slopes of the edges in $\Gamma_{0}$. The proof first shows that this solution to $M_{0}$ is unique. The final step involves continuously modifying the coefficients in $M_{0}$ to obtain a continuum of linear systems $M_{t}$, $0\leq t\leq 1$, where $M_{1}=M$. Using the fact that $M_{0}$ has a unique solution as a starting point, it is possible to show that $M_{t}$ has a unique solution for each $0\leq t\leq 1$. In particular, $M_{1}=M$ has a solution, which determines the desired straight-line crossing-free drawing of $G^{\prime}$.
@@
The proof that every free-collinear set is a free set is somewhat anticlimactic. The following argument appears in Bose et al. [17], but similar “perturb and scale” arguments are fairly common.
Let $S:=(v_{1},\ldots,v_{s})$ be a free-collinear set in $G$ and let $x_{1}<\cdots<x_{s}$ and $y_{1},\ldots,y_{s}$ be real numbers. Since $S$ is a free-collinear set in $G$, $G$ has a straight-line crossing-free drawing in $\Gamma_{0}$ where $v_{i}$ is at $(x_{i},0)$ for each $i\in\{1,\ldots,s\}$. It follows easily from the definition of straight-line crossing-free drawing that there exists some $\epsilon>0$ such that perturbing each of the vertices in $\Gamma_{0}$ by at most $\epsilon$ results in another straight-line crossing-free drawing $\Gamma_{0}^{\prime}$. Let $y:=\max\{|y_{1}|,\ldots,|y_{s}|\}$. Define $\Gamma_{0}^{\prime}$ by moving $v_{i}$ to $(x_{i},\epsilon y_{i}/y)$, for each $i\in\{1,\ldots,s\}$ and leaving the other vertices of $G$ fixed. Then each vertex in $\Gamma_{0}^{\prime}$ has been moved a distance of at most $\epsilon$ from its location in $\Gamma_{0}$, so $\Gamma_{0}^{\prime}$ is a straight-line crossing-free drawing of $G$. Finally, multiply the ${y}$-coordinate of each vertex by $y/\epsilon$ to obtain a final straight-line drawing $\Gamma$. This ${y}$-scaling is an affine transformation that does not introduce any crossings, so $\Gamma$ is a straight-line crossing-free drawing of $G$ and, for each $i\in\{1,\ldots,s\}$, $\Gamma$ places $v_{i}$ at $(x_{i},y_{i})$.
@@ 3 Graph Classes with Large Free Sets
What is the maximum size of a free set that is guaranteed to exist in any $n$-vertex planar graph? Ravsky and Verbitsky [50] observed that upper bounds on this value can be obtained from existing work on the circumference of cubic triconnected planar graphs. They observe that if a triangulation $G$ has a collinear set of size $\ell$ (and thus a large free set by 1), then its dual graph $G^{*}$ (which is a cubic triconnected planar graph) has a cycle of length $\Omega(\ell)$. The length $c(G^{*})$ of a longest cycle in $G^{*}$ is called the circumference of $G^{*}$. The circumference of cubic triconnected planar graphs has a long and rich history dating back to at least 1884 when Tait [56] conjectured that every such graph is Hamiltonian. [^11] An $n$-vertex graph is Hamiltonian if its circumference is equal to $n$.
Tait’s Conjecture was famously disproved in 1946 by Tutte, who gave an example of a non-Hamiltonian cubic triconnected planar graph having 46 vertices [57]. Repeatedly replacing vertices of Tutte’s graph with copies of itself gives a family of graphs, $\langle G_{i}:i\in\mathbb{N}\rangle$ in which $G_{i}$ has $46\cdot 45^{i}$ vertices and circumference at most $45\cdot 44^{i}$. Stated another way, $n$-vertex members of the family have circumference $O(n^{\sigma})$, for $\sigma=\log_{45}(44)<0.9941$. The current best upper bound of this type is due to Grünbaum and Walther [42] who construct a family of cubic triconnected planar graphs in which $n$-vertex members have circumference $O(n^{\sigma})$ for $\sigma=\log_{23}(22)<0.9859$. The dual of such a graph is a triangulation having $\Theta(n)$ vertices whose largest free set has size $O(n^{\sigma})$. Cano et al. [20] use this upper bound on the circumference of cubic triconnected planar graphs, along with the Erdős-Szekeres Theorem to show that $\operatorname{fix}_{\mathcal{G}}(n)\in O(n^{0.4948})$.
@@
Owens [47] constructs a family of $n$-vertex cubic triconnected planar graphs whose faces have size at most $7$, and that contain no cycle of length $\Omega(n^{0.9976})$. The dual of such a graph is a triangulation having $\Theta(n)$ vertices and maximum degree $7$ whose largest free set has size $O(n^{0.9976})$.
@@
Next, we begin by considering the subclasses of planar graphs that do admit linear-size free sets (in Section 3.1). We then study general planar graphs (in Section 3.2). The relationship between the largest free sets and the circumference is then used to obtain strong bounds for bounded-degree planar graphs (in Section 3.3).
3.1 Subclasses of planar graphs with linear-size free sets
A level planar drawing of a graph is a straight-line crossing-free drawing  in the plane, such that the vertices are placed on a sequence of horizontal lines (called levels), where each edge joins vertices in two consecutive levels. For example, a natural straight-line crossing-free drawing of a tree $T$, with root $r$, places each vertex of $T$ that is at distance $i$ from $r$ on the line $y=-i$. Thus, trees have level planar drawings.
The following easy argument shows that $n$-vertex graphs that have level planar drawings have free sets of size at least $n/2$. See Fig. 3. Starting with a level planar drawing $\Gamma$, create a simple closed curve $C$ that visits alternate levels of $\Gamma$. Either the curve that visits the even levels or the curve that visits the odd levels contains at least $\lceil n/2\rceil$ vertices of $G$. The curve $C$ does not intersect the interior of any edge of $\Gamma$. Moreover, since $C$ does not contain any two consecutive levels of $\Gamma$, the set of vertices of $G$ that are on $C$ is an independent set of $G$, so $C$ intersects each edge of $\Gamma$ in at most one endpoint. Thus $C$ is a proper-good curve.